# How Good is ChatGPT in Giving Advice on Your Visualization Design?

### v4

@@ 1. Introduction
This paper delves into ways to address this knowledge gap in practical settings. We examine the capabilities of large language models (LLMs) equipped with extensive internet-scale training data, which holds transformative potential (Okerlund et al., 2022). LLM-based chatbots, exemplified by ChatGPT (OpenAI, [n. d.]), can offer tailored suggestions by analyzing a practitioner’s design choices and considering established best practices. This emerging advancement can assist casual and novice practitioners, who may have limited knowledge and resources in data visualization, by providing human-like guidance throughout their design journey.
@@
We employ a mixed-methods approach (Fig. 1) to tackle these questions, capturing both breadth (RQ1) and depth (RQ2). RQ1 focuses on a more systematic evaluation using a broader and more realistic knowledge baseline, while RQ2 centers on the qualitative experience of receiving feedback from ChatGPT, assessed against the high standards of domain experts. In our comparisons, seasoned professionals represent trusted colleagues who bring domain expertise and rigor, whereas online practitioners provide a practical, real-world perspective, recognizing that not everyone has access to expert advice (Choi et al., 2023b). We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2023; Giannakopoulos et al., 2023; Lim et al., 2023).
@@ 2.2. Evaluating LLMs’ Knowledge Capacity
LLMs, such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020; OpenAI, 2023), and Llama (Touvron et al., 2023), have rapidly gained popularity and adoption across both industry and academia (Bommasani et al., 2021; Zhao et al., 2023). Leveraging massive amounts of text data and computational power, these foundation models have achieved impressive performance on a wide variety of tasks, from generating human-like text to high-quality images (Radford et al., 2019, 2018; Wei et al., 2022a). However, the scale and complexity of these models make it difficult to fully understand the scope and limitations of their language capabilities (Bender et al., 2021).
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2023). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
@@
Many recent studies have explored the applicability of LLMs in the realm of data visualization. Inala et al. examine the potential impact of generative AI on the data analysis process, encompassing data collection, transformation, and exploration (Inala et al., 2024). DataFormulator 1 & 2 focus on facilitating data transformation workflows for creating data visualizations (Wang et al., 2023c, 2024c). On the other hand, several tools leverage LLMs to provide tailored natural language interfaces for generating data visualizations, including LIDA (Dibia, 2023), DynaVis (Vaithilingam et al., 2024), and ChartGPT (Tian et al., 2024). Other studies have evaluated LLMs’ capabilities in eliciting design preferences (Wang et al., 2024a), assisting in chart interpretation (Wang et al., 2024b), detecting misleading charts (Lo and Qu, 2024; Alexander et al., 2024), performing visualization literacy tasks (Bendeck and Stasko, 2024; Choe et al., 2024), generating visualization code (Chen et al., 2024), and troubleshooting visualization generation code (Shen et al., 2024). Furthermore, user studies have examined how ChatGPT is utilized in classrooms (Kim et al., 2024), visualization practitioners’ perceptions of text-to-image models (Schetinger et al., 2023), the impact of AI assistance on analysis workflows (Gu et al., 2024), and its effects on visualization coding performance (Vaithilingam et al., 2022).
@@ 3.3. Analysis Decisions
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2023). That said, we discuss an automatic evaluation pipeline, outlined in the discussion section, to enable comparisons of different models without requiring human intervention.
@@ 3.4.1. Ratings on Quality Metrics
While all conditions performed well in addressing the correct question, there were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations,111visguides:visual-variables-in-vr ChatGPT 3.5 not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
@@
There were intriguing and consistent behavioral patterns across the conditions. For example, ChatGPT 3.5 responses tended to provide general, surface-level improvements applicable to any visualization, such as adding labels and legends (Figure 3). While these suggestions were often hit or miss, they were effective when the questions sought generic advice. In contrast, ChatGPT 4 exhibited a more human-like behavior, offering focused suggestions in a more authoritative manner, showcasing expertise rather than merely enumerating a list of ideas, as illustrated in Figures 3 B and C.
@@
ChatGPT 4 achieved the highest score, averaging 4.92 ($std.dev$ = 0.38), followed by ChatGPT 3.5 with an average score of 3.99 ($std.dev$ = 0.60). Human responses had the lowest average score of 3.85 ($std.dev$ = 0.94). Overall, ChatGPT 4 performed exceptionally well in providing actionable recommendations, as its responses consistently included focused, practical advice. By contrast, ChatGPT 3.5 often provided broader, less specific suggestions, which reduced its scores due to the generic or vague nature of the advice (Figures 3 B and C), although the list of recommendations often made sense (Figure 3A). The actionability scores for Human responses tend to be low when the users refrain from offering recommendations and instead validate all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset. Mann–Whitney U tests confirmed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 ($p<0.001$) and Humans ($p<0.001$). However, there was no significant difference between ChatGPT 3.5 and Human responses ($p=0.416$).
@@ 5.2. Towards automatic evaluation of LLM feedback capabilities
Facilitating ongoing monitoring of LLMs’ design feedback capabilities would be beneficial to ensure that both the development of LLMs and the systems built upon them remain well-grounded. The current reliance on human intervention to evaluate the quality of responses is inefficient. Recently, LLMs have demonstrated remarkable capabilities in evaluating natural language generation, showing greater alignment with human evaluations compared to traditional similarity-based or embedding-based metrics such as BLEU or BERTScore (Zheng et al., 2023; Liu et al., 2023b; Wang et al., 2023a; Fu et al., 2023; Kim et al., 2023).
@@
In the experiment, we employed the in-context learning method (Brown et al., 2020; Li et al., 2024) using GPT-4o-mini via API calls in December 2024, leveraging the LLM’s few-shot learning capabilities to generalize evaluation tasks from the provided examples. Based on these settings, the evaluation task was conducted in the following steps: First, the Visguides dataset was split into a few-shot development set (20%) and a test set (80%). Second, for each test case and a metric, $n$ examples were selected from the few-shot development set, where we settled on $n=5$ after testing values within a range between 1 to 8 given the limit of context lengths in the GPT-4o-mini. The process for the example selection was based on two principles: 1) Score pooling for extracting diverse examples: To ensure a diverse pool of examples ranging from low to high quality, the examples in the few-shot development set were grouped into bins for unique scores from 1 to 5 based on their scores for the target metric. 2) Embedding similarity for retrieving relevant examples: Within each bin, the most similar examples based on embedding similarity based on Sentence Transformers (Reimers and Gurevych, 2019) were selected to help guide the model to choose the examples that are most semantically close to the input query, effectively “picking the best match” from the limited example set. The number of examples was determined proportionally to the size of the corresponding bin. The selected examples were incorporated into the prompt and provided to GPT-4o-mini, released in August 2024, as presented below:
@@ 3. Specific examples from the response
We then compare the output scores to those provided by human evaluators and calculate the alignment using agreement measures and correlation coefficients. In our experiment, we use Kendall’s tau ($\tau$), Spearman ($\rho$), and Gwet’s AC1 coefficient as underlying measures. While Kendall’s tau was selected as one of the most widely used metrics, several studies (Hou, 2024; Liu et al., 2023a) have pointed out that this measure can lead to biased results when the distribution of the underlying scores is severely skewed, failing to capture the true alignment. This issue is particularly relevant in our case for metrics such as clarity, coverage, and topicality, especially for ChatGPT’s responses, where scores are skewed toward higher values (e.g., 4 or 5). To address this, Gwet’s AC1 (Gwet, 2001), which has been proposed and increasingly adopted to mitigate biased judgments, was used to complement Kendall’s tau.
Table 4 presents the outcomes of these measures for each evaluation metric. Overall, the results show that the LLM evaluator’s scores are fairly aligned with those of humans ($\tau=0.404$, $\rho=0.433$) obtained across all metrics, given a well-known interpretation of Kendall’s tau and Spearman coefficient (Akoglu, 2018; Dancey, 2007) ($\pm$0.4-0.6: moderate agreement). Specifically, the alignment was better captured for both Human and ChatGPT responses in the breadth ($\tau=0.553,\rho=0.62$; $\tau=0.522,\rho=0.637$) and actionability metric ($\tau=0.47,\rho=0.525$; $\tau=0.466,\rho=0.492$), and in the depth metric for Human responses ($\tau=0.558,\rho=0.631$). For metrics with highly skewed score distributions, such as topicality, Kendall’s tau coefficients were relatively low for ChatGPT responses. In these cases, Gwet’s AC1 coefficient provided a more nuanced perspective, demonstrating a high alignment ($\tau=-0.029,\rho=-0.029,AC1=0.915$). While our experiment shows promising capabilities of automatic evaluation using LLMs, it is worth to further investigate refining heuristics and developing prompts, for example, making prompts targeted to response types or providing clearer metric descriptions, in order to improve the alignment.
@@ 6. Design considerations for LLM-based feedback systems
It was evident that the broad knowledge base of LLMs is advantageous for rapidly exploring diverse ideas or sparking new ideas for participants. However, one issue we observed in the study was that this discovery process was mostly ad-hoc and spontaneous, resulting in a hit-or-miss experience. Future feedback systems have the potential to make this process more systematic and supportive of structured ideation. This could be achieved by providing prompts or guiding questions that steer exploration toward relevant and productive paths. For example, a recent study provided insights into a structured exploration of the design space by generating dimensions to guide response generation (Suh et al., 2023).

### v5

@@ 1. Introduction
This paper delves into ways to address this knowledge gap in practical settings. We examine the capabilities of large language models (LLMs) equipped with extensive internet-scale training data, which holds transformative potential (oke, [n. d.]). LLM-based chatbots, exemplified by ChatGPT (OpenAI, [n. d.]), can offer tailored suggestions by analyzing a practitioner’s design choices and considering established best practices. This emerging advancement can assist casual and novice practitioners, who may have limited knowledge and resources in data visualization, by providing human-like guidance throughout their design journey.
@@
We employ a mixed-methods approach (Fig. 1) to tackle these questions, capturing both breadth (RQ1) and depth (RQ2). RQ1 focuses on a more systematic evaluation using a broader and more realistic knowledge baseline, while RQ2 centers on the qualitative experience of receiving feedback from ChatGPT, assessed against the high standards of domain experts. In our comparisons, seasoned professionals represent trusted colleagues who bring domain expertise and rigor, whereas online practitioners provide a practical, real-world perspective, recognizing that not everyone has access to expert advice (Choi et al., 2023b). We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2024; Giannakopoulos et al., 2023; Lim et al., 2023).
@@ 2.2. Evaluating LLMs’ Knowledge Capacity
LLMs, such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020; OpenAI et al., 2024), and Llama (Touvron et al., 2023), have rapidly gained popularity and adoption across both industry and academia (Bommasani et al., 2021; Zhao et al., 2023). Leveraging massive amounts of text data and computational power, these foundation models have achieved impressive performance on a wide variety of tasks, from generating human-like text to high-quality images (Radford et al., 2019, 2018; Wei et al., 2022a). However, the scale and complexity of these models make it difficult to fully understand the scope and limitations of their language capabilities (Bender et al., 2021).
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2024). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
@@
Many recent studies have explored the applicability of LLMs in the realm of data visualization. Inala et al. examine the potential impact of generative AI on the data analysis process, encompassing data collection, transformation, and exploration (Inala et al., 2024). DataFormulator 1 & 2 focus on facilitating data transformation workflows for creating data visualizations (Wang et al., 2023c, 2025c). On the other hand, several tools leverage LLMs to provide tailored natural language interfaces for generating data visualizations, including LIDA (Dibia, 2023), DynaVis (Vaithilingam et al., 2024), and ChartGPT (Tian et al., 2024). Other studies have evaluated LLMs’ capabilities in eliciting design preferences (Wang et al., 2025a), assisting in chart interpretation (Wang et al., 2025b), detecting misleading charts (Lo and Qu, 2025; Alexander et al., 2024), performing visualization literacy tasks (Bendeck and Stasko, 2025; Choe et al., 2024), generating visualization code (Chen et al., 2025), and troubleshooting visualization generation code (Shen et al., 2024). Furthermore, user studies have examined how ChatGPT is utilized in classrooms (Kim et al., 2024a), visualization practitioners’ perceptions of text-to-image models (Schetinger et al., 2023), the impact of AI assistance on analysis workflows (Gu et al., 2024), and its effects on visualization coding performance (Vaithilingam et al., 2022).
@@ 3.3. Analysis Decisions
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2024). That said, we discuss an automatic evaluation pipeline, outlined in the discussion section, to enable comparisons of different models without requiring human intervention.
@@ 3.4.1. Ratings on Quality Metrics
While all conditions performed well in addressing the correct question, there were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations111visguides:visual-variables-in-vr, ChatGPT 3.5 not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
@@
There were intriguing and consistent behavioral patterns across the conditions. For example, ChatGPT 3.5 responses tended to provide general, surface-level improvements applicable to any visualization, such as adding labels and legends (Figure 3). While these suggestions were often hit or miss, they were effective when the questions sought generic advice. In contrast, ChatGPT 4 exhibited a more human-like behavior, offering focused suggestions in a more authoritative manner, showcasing expertise rather than merely enumerating a list of ideas, as illustrated in Figure 3B and C.
@@
ChatGPT 4 achieved the highest score, averaging 4.92 ($std.dev$ = 0.38), followed by ChatGPT 3.5 with an average score of 3.99 ($std.dev$ = 0.60). Human responses had the lowest average score of 3.85 ($std.dev$ = 0.94). Overall, ChatGPT 4 performed exceptionally well in providing actionable recommendations, as its responses consistently included focused, practical advice. By contrast, ChatGPT 3.5 often provided broader, less specific suggestions, which reduced its scores due to the generic or vague nature of the advice (Figure 3B and C), although the list of recommendations often made sense (Figure 3A). The actionability scores for Human responses tend to be low when the users refrain from offering recommendations and instead validate all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset. Mann–Whitney U tests confirmed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 ($p<0.001$) and Humans ($p<0.001$). However, there was no significant difference between ChatGPT 3.5 and Human responses ($p=0.416$).
@@ 5.2. Towards automatic evaluation of LLM feedback capabilities
Facilitating ongoing monitoring of LLMs’ design feedback capabilities would be beneficial to ensure that both the development of LLMs and the systems built upon them remain well-grounded. The current reliance on human intervention to evaluate the quality of responses is inefficient. Recently, LLMs have demonstrated remarkable capabilities in evaluating natural language generation, showing greater alignment with human evaluations compared to traditional similarity-based or embedding-based metrics such as BLEU or BERTScore (Zheng et al., 2023; Liu et al., 2023; Wang et al., 2023a; Fu et al., 2023; Kim et al., 2024b).
@@
In the experiment, we employed the in-context learning method (Brown et al., 2020; Li et al., 2025) using GPT-4o-mini via API calls in December 2024, leveraging the LLM’s few-shot learning capabilities to generalize evaluation tasks from the provided examples. Based on these settings, the evaluation task was conducted in the following steps: First, the Visguides dataset was split into a few-shot development set (20%) and a test set (80%). Second, for each test case and a metric, $n$ examples were selected from the few-shot development set, where we settled on $n=5$ after testing values within a range between 1 to 8 given the limit of context lengths in the GPT-4o-mini. The process for the example selection was based on two principles: 1) Score pooling for extracting diverse examples: To ensure a diverse pool of examples ranging from low to high quality, the examples in the few-shot development set were grouped into bins for unique scores from 1 to 5 based on their scores for the target metric. 2) Embedding similarity for retrieving relevant examples: Within each bin, the most similar examples based on embedding similarity based on Sentence Transformers (Devlin et al., 2018) were selected to help guide the model to choose the examples that are most semantically close to the input query, effectively “picking the best match” from the limited example set. The number of examples was determined proportionally to the size of the corresponding bin. The selected examples were incorporated into the prompt and provided to GPT-4o-mini, released in August 2024, as presented below:
@@ 3. Specific examples from the response
We then compare the output scores to those provided by human evaluators and calculate the alignment using agreement measures and correlation coefficients. In our experiment, we use Kendall’s tau ($\tau$), Spearman ($\rho$), and Gwet’s AC1 coefficient as underlying measures. While Kendall’s tau was selected as one of the most widely used metrics, several studies (Hou, 2024; Liu et al., 2023) have pointed out that this measure can lead to biased results when the distribution of the underlying scores is severely skewed, failing to capture the true alignment. This issue is particularly relevant in our case for metrics such as clarity, coverage, and topicality, especially for ChatGPT’s responses, where scores are skewed toward higher values (e.g., 4 or 5). To address this, Gwet’s AC1 (Gwet, 2001), which has been proposed and increasingly adopted to mitigate biased judgments, was used to complement Kendall’s tau.
Table 4 presents the outcomes of these measures for each evaluation metric. Overall, the results show that the LLM evaluator’s scores are fairly aligned with those of humans ($\tau=0.404$, $\rho=0.433$) obtained across all metrics, given a well-known interpretation of Kendall’s tau and Spearman coefficient (Akoglu, 2018; Dancey and Reidy, 2008) ($\pm$0.4-0.6: moderate agreement). Specifically, the alignment was better captured for both Human and ChatGPT responses in the breadth ($\tau=0.553,\rho=0.62$; $\tau=0.522,\rho=0.637$) and actionability metric ($\tau=0.47,\rho=0.525$; $\tau=0.466,\rho=0.492$), and in the depth metric for Human responses ($\tau=0.558,\rho=0.631$). For metrics with highly skewed score distributions, such as topicality, Kendall’s tau coefficients were relatively low for ChatGPT responses. In these cases, Gwet’s AC1 coefficient provided a more nuanced perspective, demonstrating a high alignment ($\tau=-0.029,\rho=-0.029,AC1=0.915$). While our experiment shows promising capabilities of automatic evaluation using LLMs, it is worth to further investigate refining heuristics and developing prompts, for example, making prompts targeted to response types or providing clearer metric descriptions, in order to improve the alignment.
@@ 6. Design considerations for LLM-based feedback systems
It was evident that the broad knowledge base of LLMs is advantageous for rapidly exploring diverse ideas or sparking new ideas for participants. However, one issue we observed in the study was that this discovery process was mostly ad-hoc and spontaneous, resulting in a hit-or-miss experience. Future feedback systems have the potential to make this process more systematic and supportive of structured ideation. This could be achieved by providing prompts or guiding questions that steer exploration toward relevant and productive paths. For example, a recent study provided insights into a structured exploration of the design space by generating dimensions to guide response generation (Suh et al., 2024).